# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ IV.3.2 Amplitude interpolation errors
As discussed in Section IV.2, we do not expect that systematic errors in our mode amplitude interpolation will significantly impact the accuracy of our waveform model. We will now test the validity of this assertion by comparing waveforms generated with two models that differ only in the generation of their amplitudes at each point output by the ODE solver: one uses amplitudes computed with our bicubic+linear interpolant (Section III.2.2), whereas the other directly solves the Teukolsky equation to obtain the mode amplitudes (Section III.2.1) using the GREMLIN code. In what follows, we will refer to waveforms from these models as “FEW” and “SAH” respectively. As the latter model is orders of magnitude more computationally expensive (despite amplitudes only being computed at $\sim 100$ points along the trajectory), we will consider three EMRI systems in lieu of a more comprehensive study over the EMRI parameter space.
@@
The first system has parameters $\{m_{1},\epsilon,a,p_{0},e_{0}\}=\{10^{6}\,M_{\odot},10^{-5},0.998,7.81,0.7\}$ (with $p_{0}$ chosen such that the CO plunges after two years), and is representative of the region of our parameter space for which our mode amplitude interpolant is least accurate (i.e., high spin and eccentricity) for inspirals of this mass ratio and duration. From Fig. 10, we do not expect weak-field amplitude interpolation errors to contribute significantly to waveform mismatch, so our choice of a two-year inspiral (over e.g., a four-year inspiral) will have little bearing on the results of this comparison. In Fig. 13, we show the cumulative mismatch (i.e., the mismatch computed over increasingly large intervals of time) between the corresponding waveforms from these two models as a function of both $t$ and $p$. Also shown in Fig. 13 is the time-domain waveform strain near the beginning and end of inspiral, with differences in mode amplitudes visible by eye in the latter case. At earlier times (in the weaker field), $\mathcal{M}\sim 10^{-8}$; this rises to $\mathcal{M}\sim 10^{-5}$ as $p\rightarrow p_{\mathrm{sep}}$. This behaviour is expected based on the poorer performance of our interpolant close to the separatrix at high spins (Fig. 10). The majority of the mismatch accumulates in the interval $p-p_{\mathrm{sep}}\in[1,0.1]$, after which the inspiral completes the few remaining orbital cycles and the cumulative mismatch levels off. To confirm that the primary cause of waveform inaccuracy near the separatrix is a result of our interpolation scheme (rather than issues of data quality), we also construct the mode amplitudes for this system with a tricubic interpolant and compare the resulting waveform with that of the “SAH” model. The cumulative mismatch we obtain (shown in the bottom panel of Fig. 13 in red) is lower than that obtained with our fiducial interpolation scheme by more than an order of magnitude, confirming that linear interpolation with respect to $a$ is indeed the limiting factor for the accuracy of our amplitude module near the separatrix.
@@
While this sharp increase in mismatch may be alarming, it is important to note that the parameters of this EMRI were chosen to probe the impact of the largest amplitude interpolation errors our model exhibits (given this mass ratio and duration). Additionally, as previously highlighted in Section II.1, the multi-scale expansion underpinning our model begins to break down in these final stages of inspiral; this region may therefore be more accurately captured by a transition-to-plunge framework Kuchler:2024esj . As mismatch only accumulates significantly near the separatrix, the deviation should be larger for more extreme mass ratios. To confirm this, we perform a similar analysis for an EMRI of mass ratio $10^{-6}$, with parameters given by the second row of Table 1. Note that $p\lesssim p_{\rm sep}+0.5$ for this entire inspiral, so the two-year duration of this observation uses wave amplitudes from the region of parameter space where our amplitude interpolation is least accurate. We do not show a visualisation of the results of this comparison for brevity, but observe a mismatch of $10^{-3}$ that steadily grows by a factor of $\sim 2$ over the course of the inspiral. As expected, this is significantly larger than what was observed for the $\epsilon=10^{-5}$ system. This highlights the inherent difficulty in modelling these EMRIs that spend many years in the neighbourhood of the separatrix; as self-force data varies rapidly in this region, it presents a significant challenge for interpolation schemes. However, for LISA, these sources will not have particularly high SNRs at typical astrophysical distances — the SNR of this source is $30$ — meaning that (according to Eq. 15) even given the poorer performance of our model in this region of the parameter space, we still do not expect to observe signficiant systematic biases during parameter estimation as a result. While good performance over the entire parameter space is obviously desirable, there is also scope in future work to optimise waveform models based on the expected SNRs of astrophysical sources, which may improve the efficiency of these models without introducing significant biases into the results of data analysis.
@@ V Results and discussion
In previous sections, we have described the implementation of our waveform model and demonstrated it is robust over its domain of validity. Here, we apply our model to explore scientific and data analysis prospects for asymmetric-mass binary observations with LISA. We first characterise the computational cost of our waveform model (Section V.1), which is closely tied to the feasibility of data analysis and parameter estimation. We then examine the impact of semi-relativistic models for waveform mode amplitudes (which have been applied extensively in the literature) on the SNR of EMRI sources (Section V.2), and explore how these approximate models can induce biases during parameter recovery in Section V.3. Finding that SNRs computed with approximate models are subject to significant errors, in Section V.4 we apply our model to the exploration of the sky-averaged detection horizon for eccentric and rapidly-spinning EMRI and IMRI systems, examining its dependence on the component masses of the system. We then examine the detectability of small eccentricities for a representative, rapidly-spinning EMRI system in Section V.5, which is an area of great astrophysical interest with respect to the formation of EMRIs in gas-dominated environments such as accretion disks. We conclude this section with an exploration of parameter recovery for EMRI and IMRI sources (with parameters listed in  Table 1) in the Bayesian inference context, highlighting similarities and differences in the results obtained between these two cases.
@@ V.1 Waveform computational cost analysis
The principal objective of the few package is to provide tools for the rapid generation of adiabatic waveforms by leveraging vectorised operations on GPU hardware. In this subsection, we will examine the computational cost of waveform generation in both the time domain (TD) and the frequency domain (FD). The procedures for constructing waveforms in these two domains differ only at the stage of the waveform summation modules, which have not changed significantly since the TD (Refs. Katz:2021yft ; chua2021rapid ) and FD (Ref. Speri:2023jte ) releases of the package. The differences in timing with respect to these previous works can therefore be attributed to: the expansion of the parameter-space coverage of the model, particularly the inclusion of spin and the evolution of the inspiral closer to the separatrix; and changes made to the trajectory and amplitude generation modules. We warn the reader that the computational cost of the waveform can vary depending on the computing resources used; all computational wall-times reported in this section were obtained using an NVIDIA A100 GPU and a $2\,\mathrm{GHz}$ AMD EPYC-7713 CPU. Furthermore, the timing distribution will depend on our procedure for sampling source parameters — this is described in Appendix C.
@@
We will examine the cost of our model for four-year inspirals with sampling cadence $\mathrm{d}t=5\,\mathrm{s}$. We consider two mode selection thresholds, $\kappa=(10^{-2},10^{-5})$, to probe how this cost depends on the mode content requested in waveform generation. Drawing $10^{4}$ sets of parameters, we show the corresponding distribution of computational wall-times per waveform evaluation in Fig. 14. We obtain median wall-times of $0.11\,\mathrm{s}$ ($0.13\,\mathrm{s}$) per TD evaluation and $0.13\,\mathrm{s}$ ($0.15\,\mathrm{s}$) per FD evaluation, for $\kappa=10^{-2}\,(10^{-5})$. The FD wall-time distribution presents a long upper tail that extends to $\sim 1\,\mathrm{s}$; this feature corresponds to primary masses $m_{1}\lesssim 10^{6}\,M_{\odot}$ with $e_{0}\gtrsim 0.7$. The reason for this discrepancy in timings between output domains is as follows (based on arguments presented in Ref. Speri:2023jte ). The computational cost of the waveform scales with the number of data points at which each waveform mode must be evaluated. For TD generation, each array element receives contributions from all the selected harmonics (such that the overall summation cost scales linearly with the number of waveform modes). In contrast, the cost-per-mode of the FD generation depends on the frequency evolution of that mode, as each mode is only evaluated in frequency bins it evolves through. Therefore, the FD waveform is most efficient when the modes of the signal span a compact range of frequencies, which corresponds to higher primary mass values; conversely, at lower primary masses and higher eccentricities many modes span a wide range of frequencies, leading to the tail observed in Fig. 14. In terms of performance with respect to eccentricity at higher primary masses, it is argued in Speri:2023jte that FD waveform generation is more efficient than its TD counterpart for $e_{0}\sim 0.7$ when $m_{1}\sim 10^{7}$, and this remains true in our results.
The inclusion of primary spin increases the azimuthal frequency $\hat{\Omega}_{\phi}$ close to the separatrix by up to a factor of six (for $a=0.999,e=0)$, which in turn leads to many modes evolving over a much larger range of frequencies in a typical waveform than for the zero-spin case. This causes the FD waveform generation to exceed that of TD generation for the majority of spins $a>0$ (which is why the lower ends of the FD timing distributions do not extend below their TD counterparts). We remind the reader that Fig. 14 does not include retrograde orbits ($a<0$), where we expect the opposite behaviour to occur and the relative cost of FD waveform generation to decrease.
@@ V.2 Impact of relativistic amplitudes on waveform SNR
Semi-relativistic waveform models known as “kludges” barack2004lisa ; babak2007kludge ; chua2017augmented have been commonly-used tools to make qualitative statements describing detection rates, to perform large-scale parameter measurement studies, and to examine the impact of stochastic backgrounds composed of unresolved EMRI signals on LISA data analysis Babak:2017tow ; bonetti2020gravitational ; piarulli2024test ; Pozzoli:2023kxy ; Chapman-Bird:2022tvu . These studies perform SNR calculations with kludge models that only incorporate a quadrupolar approximation (i.e., $\ell=2=m$) to the mode structure of the waveform. As was identified in Refs. Fujita:2020zxe ; Khalvati:2024tzz , these approximate quadrupolar amplitudes may be insufficiently accurate to adequately describe the EMRI mode spectrum, particularly in the strong-field regime of gravity accessible to eccentric inspirals into rapidly-spinning black holes. As we have already shown in Fig. 4, a large number of $(\ell,m,n)$ modes are required to faithfully reconstruct the true EMRI signal. In this subsection, we will investigate through SNR comparisons (which are critical for determining detection rates) the impact of approximating the mode amplitudes of adiabatic EMRI waveforms with those of the Augmented Analytic Kludge (AAK), a kludge model that has seen extensive application in the literature. Understanding the regions of parameter space where the SNR is inaccurately estimated, and to what extent, is crucial for avoiding misrepresentations of the detection rate of EMRIs.
@@
Exploiting the modular nature of few, we combine our relativistic trajectory model (Section III.1) with two different amplitude models. This results in two waveform models: the first uses our adiabatic mode amplitudes (Section III.2) and is the model we introduce in this work; the second uses the AAK mode amplitudes. We will refer to these models as Kerr and AAK respectively throughout this subsection. The Kerr amplitudes, described in  Section III.2, consider the harmonics $(\ell,m,n)$ for $2\leq\ell\leq 10$, $|m|\leq\ell$, and $|n|\leq 55$. For the AAK model, which is quadrupolar Blanchet:2013haa (i.e., $\ell=2=m$), we set the number of Fourier modes to $n_{\mathrm{max}}=50$; note that these modes do not have a one-to-one correspondence with those of our model (see Ref. chua2017augmented for definitions). We confirmed that this value of $n_{\mathrm{max}}$ is sufficiently high to represent the AAK for the purposes of this analysis (i.e., our results are convergent with respect to $n_{\mathrm{max}}$). For the Kerr model, we use the default mode-selection threshold $\kappa=10^{-5}$; by construction, the SNRs computed with the Kerr model for $\kappa=0$ and $\kappa=10^{-5}$ will differ by $\sim 10^{-5}$, which is negligible on the scale of the SNR variations we identify in our results.
@@
In our analysis, we adopt the data analysis conventions and methods outlined in Appendix D. For each set of source parameters, we choose $p_{0}$ such that inspirals plunge after two years and fix extrinsic parameters (such as orientation angles and initial phases) to those given in the caption of Table 1. Our results are largely independent of this choice of extrinsic parameters.
@@
Figure 15: Ratio of the optimal SNRs computed with the semi-relativistic AAK mode amplitude model ($\rho_{\mathrm{AAK}}$) and the adiabatic model presented in this work ($\rho_{\mathrm{Kerr}}$), as a function of initial eccentricity $e_{0}\in[0,0.7]$ and primary spin $a\in[0,0.998]$. For all systems, we consider four-year inspirals with masses $(m_{1},m_{2})=(10^{6},10)\,M_{\odot}$. Waveforms for both models are built with the same adiabatic trajectory model (Section III.1). The (dashed, solid, dot-dashed) white lines indicate the ratio of the SNRs for three reference values $\rho_{\rm AAK}/\rho_{\rm Kerr}\in\{1.1,1,0.9\}$.
@@
In Fig. 15, we plot the SNR $\rho$ as a function of $e_{0}$ and $a$ for EMRIs with $(m_{1},m_{2})=(10^{6},10)\,M_{\odot}$. The solid white line indicates the boundary where the two waveform models provide a consistent prediction of the signal-to-noise ratio $\rho_{\text{AAK}}/\rho_{\text{Kerr}}=1$. Above this line (at higher spins), the AAK overestimates the SNR by up to $\sim$30%, whereas below it underestimates it by up to $\sim$20%. This discrepancy between the models increases with eccentricity, particularly at lower values of $a$. Two main factors contribute to the differences observed. First, as the AAK is quadrupolar, it will become increasingly inaccurate as the orbital separation decreases and missing modes with $\ell>2$ become more significant. Second, as the mode amplitudes of the AAK are based upon a low-velocity approximation Peters:1963ux , they will also become increasingly inaccurate as the orbital separation decreases. What we observe in Fig. 15 is a combination of these two effects.

### v2

@@ IV.3.2 Amplitude interpolation errors
As discussed in Section˜IV.2, we do not expect that systematic errors in our mode amplitude interpolation will significantly impact the accuracy of our waveform model. We will now test the validity of this assertion by comparing waveforms generated with two models that differ only in the generation of their amplitudes at each point output by the ODE solver: one uses amplitudes computed with our bicubic+linear interpolant (Section˜III.2.2), whereas the other directly solves the Teukolsky equation to obtain the mode amplitudes (Section˜III.2.1) using the GREMLIN code. In what follows, we will refer to waveforms from these models as “FEW” and “SAH” respectively. As the latter model is orders of magnitude more computationally expensive (despite amplitudes only being computed at $\sim 100$ points along the trajectory), we will consider three EMRI systems in lieu of a more comprehensive study over the EMRI parameter space.
@@
The first system has parameters $\{m_{1},\epsilon,a,p_{0},e_{0}\}=\{10^{6}\,M_{\odot},10^{-5},0.998,7.81,0.7\}$ (with $p_{0}$ chosen such that the CO plunges after two years), and is representative of the region of our parameter space for which our mode amplitude interpolant is least accurate (i.e., high spin and eccentricity) for inspirals of this mass ratio and duration. From Fig.˜10, we do not expect weak-field amplitude interpolation errors to contribute significantly to waveform mismatch, so our choice of a two-year inspiral (over e.g., a four-year inspiral) will have little bearing on the results of this comparison. In Fig.˜13, we show the cumulative mismatch (i.e., the mismatch computed over increasingly large intervals of time) between the corresponding waveforms from these two models as a function of both $t$ and $p$. Also shown in Fig.˜13 is the time-domain waveform strain near the beginning and end of inspiral, with differences in mode amplitudes visible by eye in the latter case. At earlier times (in the weaker field), $\mathcal{M}\sim 10^{-8}$; this rises to $\mathcal{M}\sim 10^{-5}$ as $p\rightarrow p_{\mathrm{sep}}$. This behaviour is expected based on the poorer performance of our interpolant close to the separatrix at high spins (Fig.˜10). The majority of the mismatch accumulates in the interval $p-p_{\mathrm{sep}}\in[1,0.1]$, after which the inspiral completes the few remaining orbital cycles and the cumulative mismatch levels off. To confirm that the primary cause of waveform inaccuracy near the separatrix is a result of our interpolation scheme (rather than issues of data quality), we also construct the mode amplitudes for this system with a tricubic interpolant and compare the resulting waveform with that of the “SAH” model. The cumulative mismatch we obtain (shown in the bottom panel of Fig.˜13 in red) is lower than that obtained with our fiducial interpolation scheme by more than an order of magnitude, confirming that linear interpolation with respect to $a$ is indeed the limiting factor for the accuracy of our amplitude module near the separatrix.
@@
While this sharp increase in mismatch may be alarming, it is important to note that the parameters of this EMRI were chosen to probe the impact of the largest amplitude interpolation errors our model exhibits (given this mass ratio and duration). Additionally, as previously highlighted in Section˜II.1, the multi-scale expansion underpinning our model begins to break down in these final stages of inspiral; this region may therefore be more accurately captured by a transition-to-plunge framework Kuchler:2024esj . As mismatch only accumulates significantly near the separatrix, the deviation should be larger for more extreme mass ratios. To confirm this, we perform a similar analysis for an EMRI of mass ratio $10^{-6}$, with parameters given by the second row of Table˜1. Note that $p\lesssim p_{\rm sep}+0.5$ for this entire inspiral, so the two-year duration of this observation uses wave amplitudes from the region of parameter space where our amplitude interpolation is least accurate. We do not show a visualisation of the results of this comparison for brevity, but observe a mismatch of $10^{-3}$ that steadily grows by a factor of $\sim 2$ over the course of the inspiral. As expected, this is significantly larger than what was observed for the $\epsilon=10^{-5}$ system. This highlights the inherent difficulty in modelling these EMRIs that spend many years in the neighbourhood of the separatrix; as self-force data varies rapidly in this region, it presents a significant challenge for interpolation schemes. However, for LISA, these sources will not have particularly high SNRs at typical astrophysical distances — the SNR of this source is $30$ — meaning that (according to Eq.˜15) even given the poorer performance of our model in this region of the parameter space, we still do not expect to observe signficiant systematic biases during parameter estimation as a result. While good performance over the entire parameter space is obviously desirable, there is also scope in future work to optimise waveform models based on the expected SNRs of astrophysical sources, which may improve the efficiency of these models without introducing significant biases into the results of data analysis.
@@ V Results and discussion
In previous sections, we have described the implementation of our waveform model and demonstrated it is robust over its domain of validity. Here, we apply our model to explore scientific and data analysis prospects for asymmetric-mass binary observations with LISA. We first characterise the computational cost of our waveform model (Section˜V.1), which is closely tied to the feasibility of data analysis and parameter estimation. We then examine the impact of semi-relativistic models for waveform mode amplitudes (which have been applied extensively in the literature) on the SNR of EMRI sources (Section˜V.2), and explore how these approximate models can induce biases during parameter recovery in Section˜V.3. Finding that SNRs computed with approximate models are subject to significant errors, in Section˜V.4 we apply our model to the exploration of the sky-averaged detection horizon for eccentric and rapidly-spinning EMRI and IMRI systems, examining its dependence on the component masses of the system. We then examine the detectability of small eccentricities for a representative, rapidly-spinning EMRI system in Section˜V.5, which is an area of great astrophysical interest with respect to the formation of EMRIs in gas-dominated environments such as accretion disks. We conclude this section with an exploration of parameter recovery for EMRI and IMRI sources (with parameters listed in Table˜1) in the Bayesian inference context in Section˜V.6, highlighting similarities and differences in the results obtained between these two cases. Unless otherwise stated, our analysis procedures throughout this section follow the conventions outlined in Appendix˜D.
@@ V.1 Waveform computational cost analysis
The principal objective of the few package is to provide tools for the rapid generation of adiabatic waveforms by leveraging vectorised operations on GPU hardware. In this subsection, we will examine the computational cost of waveform generation in both the time domain (TD) and the frequency domain (FD). The procedures for constructing waveforms in these two domains differ only at the stage of the waveform summation modules, which have not changed significantly since the TD (Refs. Katz:2021yft ; chua2021rapid ) and FD (Ref. Speri:2023jte ) releases of the package. The differences in timing with respect to these previous works can therefore be attributed to: the expansion of the parameter-space coverage of the model, particularly the inclusion of spin and the evolution of the inspiral closer to the separatrix; and changes made to the trajectory and amplitude generation modules. We warn the reader that the computational cost of the waveform can vary depending on the computing resources used; all computational wall-times reported in this section were obtained using an NVIDIA A100 GPU and a $2\,\mathrm{GHz}$ AMD EPYC-7713 CPU. Furthermore, the timing distribution will depend on our procedure for sampling source parameters — this is described in Appendix˜C.
@@
We will examine the cost of our model for four-year inspirals with sampling cadence $\mathrm{d}t=5\,\mathrm{s}$. This sampling cadence is based roughly on what is required for systems with $M\sim 10^{6}\,M_{\odot}$; the computational cost for other choices of cadence can be roughly estimated by scaling the wall-time with respect to this value of $\mathrm{d}t$. We consider two mode selection thresholds, $\kappa=(10^{-2},10^{-5})$, to probe how this cost depends on the mode content requested in waveform generation. Drawing $10^{4}$ sets of parameters, we show the corresponding distribution of computational wall-times per waveform evaluation in Fig.˜14. We obtain median wall-times of $0.11\,\mathrm{s}$ ($0.13\,\mathrm{s}$) per TD evaluation and $0.13\,\mathrm{s}$ ($0.15\,\mathrm{s}$) per FD evaluation, for $\kappa=10^{-2}\,(10^{-5})$. The FD wall-time distribution presents a long upper tail that extends to $\sim 1\,\mathrm{s}$; this feature corresponds to primary masses $m_{1}\lesssim 10^{6}\,M_{\odot}$ with $e_{0}\gtrsim 0.7$. The reason for this discrepancy in timings between output domains is as follows (based on arguments presented in Ref. Speri:2023jte ). The computational cost of the waveform scales with the number of data points at which each waveform mode must be evaluated. For TD generation, each array element receives contributions from all the selected harmonics (such that the overall summation cost scales linearly with the number of waveform modes). In contrast, the cost-per-mode of the FD generation depends on the frequency evolution of that mode, as each mode is only evaluated in frequency bins it evolves through. Therefore, the FD waveform is most efficient when the modes of the signal span a compact range of frequencies, which corresponds to higher primary mass values; conversely, at lower primary masses and higher eccentricities many modes span a wide range of frequencies, leading to the tail observed in Fig.˜14. In terms of performance with respect to eccentricity at higher primary masses, it is argued in Speri:2023jte that FD waveform generation is more efficient than its TD counterpart for $e_{0}\sim 0.7$ when $m_{1}\sim 10^{7}$, and this remains true in our results. It is worth noting that frequency-domain waveforms generated via the stationary phase approximation can be downsampled to reduce computational cost, potentially improving efficiency for certain analyses Speri:2023jte . However, such downsampling reduces the effective signal-to-noise ratio, leading to biased parameter posteriors in noisy data, limiting its applicability for realistic EMRI inference.
The inclusion of primary spin increases the azimuthal frequency $\hat{\Omega}_{\phi}$ close to the separatrix by up to a factor of six (for $a=0.999,e=0)$, which in turn leads to many modes evolving over a much larger range of frequencies in a typical waveform than for the zero-spin case. This causes the FD waveform generation to exceed that of TD generation for the majority of spins $a>0$ (which is why the lower ends of the FD timing distributions do not extend below their TD counterparts). We remind the reader that Fig.˜14 does not include retrograde orbits ($a<0$), where we expect the opposite behaviour to occur and the relative cost of FD waveform generation to decrease.
@@ V.2 Impact of relativistic amplitudes on waveform SNR
Semi-relativistic waveform models known as “kludges” barack2004lisa ; babak2007kludge ; chua2017augmented have been commonly-used tools to make qualitative statements describing detection rates, to perform large-scale parameter measurement studies, and to examine the impact of stochastic backgrounds composed of unresolved EMRI signals on LISA data analysis Babak:2017tow ; bonetti2020gravitational ; piarulli2024test ; Pozzoli:2023kxy ; Chapman-Bird:2022tvu . These studies perform SNR calculations with kludge models that only incorporate a quadrupolar approximation (i.e., $\ell=2=m$) to the mode structure of the waveform. As was identified in Refs. Fujita:2020zxe ; Khalvati:2024tzz , these approximate quadrupolar amplitudes may be insufficiently accurate to adequately describe the EMRI mode spectrum, particularly in the strong-field regime of gravity accessible to eccentric inspirals into rapidly-spinning black holes. As we have already shown in Fig.˜4, a large number of $(\ell,m,n)$ modes are required to faithfully reconstruct the true EMRI signal. In this subsection, we will investigate through SNR comparisons (which are critical for determining detection rates) the impact of approximating the mode amplitudes of adiabatic EMRI waveforms with those of the Augmented Analytic Kludge (AAK), a kludge model that has seen extensive application in the literature. Understanding the regions of parameter space where the SNR is inaccurately estimated, and to what extent, is crucial for avoiding misrepresentations of the detection rate of EMRIs.
@@
Exploiting the modular nature of few, we combine our relativistic trajectory model (Section˜III.1) with two different amplitude models. This results in two waveform models: the first uses our adiabatic mode amplitudes (Section˜III.2) and is the model we introduce in this work; the second uses the AAK mode amplitudes. We will refer to these models as Kerr and AAK respectively throughout this subsection. The Kerr amplitudes, described in Section˜III.2, consider the harmonics $(\ell,m,n)$ for $2\leq\ell\leq 10$, $|m|\leq\ell$, and $|n|\leq 55$. For the AAK model, which is quadrupolar Blanchet:2013haa (i.e., $\ell=2=m$), we set the number of Fourier modes to $n_{\mathrm{max}}=50$; note that these modes do not have a one-to-one correspondence with those of our model (see Ref. chua2017augmented for definitions). We confirmed that this value of $n_{\mathrm{max}}$ is sufficiently high to represent the AAK for the purposes of this analysis (i.e., our results are convergent with respect to $n_{\mathrm{max}}$). For the Kerr model, we use the default mode-selection threshold $\kappa=10^{-5}$; by construction, the SNRs computed with the Kerr model for $\kappa=0$ and $\kappa=10^{-5}$ will differ by $\sim 10^{-5}$, which is negligible on the scale of the SNR variations we identify in our results.
@@
For each set of source parameters, we choose $p_{0}$ such that inspirals plunge after four years and fix extrinsic parameters (such as orientation angles and initial phases) to those given in the caption of Table˜1. Our results are largely independent of this choice of extrinsic parameters.
@@
Figure 15: Ratio of the optimal SNRs computed with the semi-relativistic AAK mode amplitude model ($\rho_{\mathrm{AAK}}$) and the adiabatic model presented in this work ($\rho_{\mathrm{Kerr}}$), as a function of initial eccentricity $e_{0}\in[0,0.7]$ and primary spin $a\in[0,0.998]$. For all systems, we consider four-year inspirals with masses $(m_{1},m_{2})=(10^{6},10)\,M_{\odot}$. Waveforms for both models are built with the same adiabatic trajectory model (Section˜III.1). The (dashed, solid, dot-dashed) white lines indicate the ratio of the SNRs for three reference values $\rho_{\rm AAK}/\rho_{\rm Kerr}\in\{1.1,1,0.9\}$.
@@
In Fig.˜15, we plot the SNR $\rho$ as a function of $e_{0}$ and $a$ for EMRIs with $(m_{1},m_{2})=(10^{6},10)\,M_{\odot}$. The solid white line indicates the boundary where the two waveform models provide a consistent prediction of the signal-to-noise ratio $\rho_{\text{AAK}}/\rho_{\text{Kerr}}=1$. Above this line (at higher spins), the AAK overestimates the SNR by up to $\sim$30%, whereas below it underestimates it by up to $\sim$20%. This discrepancy between the models increases with eccentricity, particularly at lower values of $a$. Two main factors contribute to the differences observed. First, as the AAK is quadrupolar, it will become increasingly inaccurate as the orbital separation decreases and missing modes with $\ell>2$ become more significant. Second, as the mode amplitudes of the AAK are based upon a low-velocity approximation Peters:1963ux , they will also become increasingly inaccurate as the orbital separation decreases. What we observe in Fig.˜15 is a combination of these two effects.